# The Devil behind the mask: An emergent safety vulnerability of Diffusion LLMs

### v1

@@ 1 Introduction
Diffusion-based language models (dLLMs) (Ye et al., 2025; Nie et al., 2025b) have recently emerged as a promising complementary paradigm to traditional autoregressive LLMs (Achiam et al., 2023; Yang et al., 2024). Unlike sequential generation, dLLMs support parallel decoding of masked tokens and leverage bidirectional context modeling, enabling theoretically faster inference and more holistic understanding of input prompts (Yu et al., 2025). These advantages have led to impressive performance and efficiency in tasks such as code generation (Labs et al., 2025; Gong et al., 2025), complex reasoning (Zhu et al., 2025), and text infilling (Li et al., 2025). Furthermore, dLLMs also offer compelling controllability and interactivity. Specifically, users can flexibly insert masked tokens at arbitrary positions in the instruction or generated content, allowing for precise, context-aware editing or rewriting, formatted generation, and structured information extraction as shown in Figure 1.
@@
Although the efficiency and interactivity enabled by parallel decoding and bidirectional context modeling highlight the great potential and promising applications of dLLMs, they may also expose new vulnerabilities in safety. Specifically, dLLMs depart from the left-to-right, causal generation paradigm (Achiam et al., 2023; Touvron et al., 2023) and instead leverage bidirectional context modeling. At each denoising step, the dLLMs can “see” the entire token sequence around every [MASK] and fill in tokens that best preserve overall coherence and fluency. While this design yields remarkable flexibility, inserting or regenerating arbitrary spans without retraining, it also creates a blind spot in current safety alignment techniques. By interleaving sets of [MASK] tokens after vanilla malicious prompt, as shown in Figure 2, a dLLM is coerced into generating harmful instructions purely to maintain contextual consistency.
@@
Moreover, in contrast to autoregressive LLMs, which generate tokens sequentially and can perform on-the-fly rejection of unsafe continuations, dLLMs decode masked tokens in parallel at each step, substantially limiting the model’s ability to conduct dynamic risk assessment or intervene during generation (e.g., reject sampling for tokens corresponding to harmful contents). Consequently, defenses designed for left-to-right models break down, opening the door to powerful new jailbreak attacks.
@@
Our contributions are summarized as follows:
@@ 2.1 Diffusion Large Language Models
Diffusion Models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have significantly advanced the field of generative modeling, particularly in continuous domains such as images (Rombach et al., 2022; Peebles & Xie, 2023). However, extending these models to discrete data like text introduces distinct challenges due to the inherent discreteness of language. A promising direction in this space is the development of Masked Diffusion Models (MDMs) (Austin et al., 2021; Lou et al., 2023; Shi et al., 2024; Nie et al., 2025a; b; Hoogeboom et al., 2021; Campbell et al., 2022), which generate text by iteratively predicting masked tokens conditioned on their surrounding context. This approach has emerged as a compelling alternative to the traditional autoregressive framework in large language models (LLMs), opening new avenues for text generation. Noteworthy instances of MDMs include LLaDA (Nie et al., 2025b), an 8-billion-parameter model trained from scratch with a bidirectional Transformer architecture, and Dream (Ye et al., 2025), which builds upon pre-trained autoregressive model (ARM) weights. Both models achieve performance comparable to similarly sized ARMs such as LLaMA3 8B (Dubey et al., 2024). The bidirectional nature of these models offers potential advantages over ARMs, including mitigating issues like the reversal curse (Berglund et al., 2023), thus positioning diffusion-based methods as a competitive alternative for next-generation foundation language models.
2.2 Jailbreak Attacks and Defenses
@@
Recent studies reveal diverse jailbreak attacks on LLMs by treating them as either computational systems or cooperative agents (Ren et al., 2024; Chen et al., 2024c). Search-based methods like GCG (Zou et al., 2023), AutoDAN (Liu et al., 2023), and PAIR (Chao et al., 2025) use optimization or genetic algorithms to generate adversarial prompts, while side-channel attacks exploit low-resource languages to evade safety checks (Deng et al., 2023). Other techniques target LLMs’ weaknesses in reasoning and symbolic understanding, including scenario nesting (Ding et al., 2023), prompt decomposition (Li et al., 2024), and ASCII obfuscation (Jiang et al., 2024). Additionally, some attacks anthropomorphize LLMs, inducing harmful outputs through persuasion or cognitive overload (Li et al., 2023; Zeng et al., 2024; Xu et al., 2023). To mitigate these threats, defenses fall into four main categories: (1) filter-based detection via perplexity or external classifiers (Jain et al., 2023; Phute et al., 2023; Chen et al., 2024b; ); (2) input modification through permutation or paraphrasing (Robey et al., 2023); (3) prompt-based reminders to reinforce ethical behavior (Xie et al., 2023); and (4) optimization-based approaches such as robust prompt design or safe alignment (Qian et al., 2024; Zhou et al., 2024; Xu et al., 2024; Lu et al., 2025). However, these methods are predominantly developed for autoregressive LLMs. The jailbreak safety of diffusion-based LLMs remains largely unexplored, leaving an important open problem for future research.
@@ 3.1 Preliminary
[…]
@@
$P_{\phi}(\mathbf{y}|\mathbf{c},\mathbf{y}^{(k)})=f_{\phi}(\mathbf{c},\mathbf{y}^{(k)};\phi).$ (2)
@@
[…]
$\mathbf{x}^{(k-1)}=S\left(\operatorname*{arg\,max}_{\mathbf{y}}P_{\phi}(\mathbf{y}|\mathbf{c},\mathbf{x}^{(k)}),\mathbf{x}^{(k)},\mathbf{c},k\right),$ (5)
where $\mathbf{x}^{(K)}$ initializes with new mask tokens inserted at user-specified positions. This facilitates:
  • Targeted regeneration by masking unsatisfactory spans $\mathbf{x}_{[i:j]}$.
@@ 3.2 DiJA: Diffusion-based LLMs Jailbreak Attack
3.2.1 Motivation
We begin by constructing the corresponding interleaved mask-text jailbreak prompt based on the original jailbreak prompt (e.g., harmful behaviors from Harmbench (Mazeika et al., 2024)). Let $\mathbf{a}=(a_{1},...,a_{R})$ be a jailbreak prompt and $\mathbf{m}=(\texttt{[MASK]},...,\texttt{[MASK]})_{Q}$ be $Q$ consecutive masks. An interleaved mask-text jailbreak prompt can be constructed:
@@ 3.2.1 Motivation
where $\oplus$ denotes concatenation, $\otimes$ interleaving, and $\mathbf{w}$ benign separator text. It is worth noting that our constructed prompt does not obscure or remove any of the hazardous content present in the vanilla jailbreak prompt. This interleaved mask-text prompt construction enables forced generation at specific masked positions, which fundamentally bypasses alignment safeguards in dLLMs. Formally, given an interleaved mask-text prompt $\mathbf{p}_{\text{i}}$, the model’s output distribution factorizes as:
$P_{\phi}(\mathbf{y}|\mathbf{p}_{\text{i}})=\prod{t\in\mathcal{M}}P_{\phi}(y_{t}|{\mathbf{p}_{\text{i}}\setminus t})\cdot\prod{t\notin\mathcal{M}}\delta(y_{t}=p_{t}),$ (7)
@@
3.2.2 Method Design
Specifically, we leverage a language model (e.g., Qwen2.5-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct or GPT-4o) to automatically construct interleaved mask-text jailbreak prompts via in-context learning. The in-context learning template can be found in Appendix A.5. To ensure the generalization and effectiveness of DiJA in jailbreak attacking, we introduce three strategies into the in-context learning process, aiming to enhance the diversity and coherence of the constructed interleaved mask-text jailbreak prompts.
Prompt Diversification. To ensure the diversity of interleaved mask-text jailbreak prompts, it is essential to first guarantee the diversity of the underlying vanilla jailbreak prompts from which they are constructed. We manually curate a small yet diverse set of jailbreak examples as few-shot demonstrations for in-context learning. These examples span a variety of jailbreak prompt forms (e.g., step-by-step guides, Q&A, lists, markdowns, dialogues, emails) and harmful content (e.g., malware generation, phishing schemes, hate speech, illegal drug recipes, violence instructions ), ensuring robustness against surface-level prompt variations. We further inject stylistic perturbations (e.g., tone, formality, verbosity) to simulate realistic adversarial scenarios and prevent overfitting.
Masking Pattern Selection. Building on the diversified vanilla prompts, we apply a range of masking strategies to further enhance the diversity of masking patterns. These include: Block-wise masking, which masks entire spans to simulate redacted instructions and elicit long, coherent generations; Fine-grained masking, which selectively hides key tokens (e.g., verbs or entities) while preserving structure; and Progressive masking, which incrementally masks critical information across multi-step instructions to amplify intent. Each strategy balances contextual anchoring with generative freedom, allowing fine-grained control over dLLM behavior and broadening attack coverage. Illustrative examples are provided in Table 6 (Appendix A.5).
Benign Separator Insertion. After ensuring diversity in both the content and structure of the vanilla prompts as well as in the masking patterns, a crucial step lies in preserving the fluency and coherence of the final interleaved mask-text prompts. This involves carefully aligning the vanilla prompt segments with the masked tokens to maximize the effectiveness of the attack. To this end, we insert short, harmless snippets drawn from a curated phrase pool or generated via controlled prompts. These separators are stylistically consistent (e.g., factual, instructive, narrative), semantically neutral, and capped at ten words. They serve two key purposes: (i) preserving fluency and structural coherence, and (ii) anchoring context to guide dLLMs toward harmful completions. Importantly, the separators are context-sensitive—adapted to the rhetorical style of the original jailbreak prompt (e.g., procedural, persuasive, or conversational)—to ensure seamless integration and stealth. This alignment helps the model treat masked spans as natural continuations, improving attack success without sacrificing realism.
@@ 3.2.2 Method Design
1:Vanilla jailbreak prompt $\mathbf{a}=(a_{1},\dots,a_{R})$
@@
3:Examples of interleaved text-mask prompts =$\mathcal{E}=\{(\mathbf{a}^{(i)},\mathbf{p}^{(i)}_{\text{i}})\}_{i=1}^{K}$
4:Refinement language model $\mathcal{L}$; target diffusion-based LLM $\mathcal{D}$
@@
6:// Prompt Construction & Refinement
@@
8:Provide few-shot examples of interleaved prompts and vanilla jailbreak prompt $\mathbf{a}$ to $\mathcal{L}$
@@
11:// Masked Decoding in Diffusion-based LLMs
@@
14:     Sample token: $y_{t}\sim P_{\phi}(y_{t}\mid\mathbf{p}_{\text{i}}\setminus t)$triangleright The decoding of [MASK] is performed in parallel
@@ 4.1 Experiment Settings
Implementation Details. To evaluate the effectiveness of our proposed automated jailbreak attack pipeline and uncover critical security vulnerabilities in existing diffusion-based LLMs (dLLMs), we conduct experiments on representative dLLMs, including the LLaDA family (Nie et al., 2025b), Dream family (Ye et al., 2025), and MMaDA family (Yang et al., 2025), across multiple recognized jailbreak benchmarks (Mazeika et al., 2024; Chao et al., 2024a; Souly et al., 2024). We experimented with two LLMs for constructing and refining interleaved mask-text jailbreak prompts in DiJA: Qwen2.5-7B-Instruct (Yang et al., 2024) (denoted as DiJA) and GPT-4o (Hurst et al., 2024) (denoted as DiJA*), with results reported in Tables 12, and 3, respectively. For more details on the victim models, benchmarks, and baselines, please refer to the Appendix A.
Evaluation Metrics. Building on prior works (Liu et al., 2023; Chao et al., 2023; Ding et al., 2023; Dong et al., 2024; Chen et al., 2024d), we evaluate using GPT-judged Harmful Score (HS) and Attack Success Rate (ASR), including keyword-based ASR (ASR-k) and evaluator-based ASR (ASR-e). GPT-4o rates victim model responses from 1 (refusal or harmless) to 5 (highly harmful or relevant), HS=5 denotes a successful jailbreak. We use the same judging prompt as in previous studies. For more details on the evaluation metrics, please refer to the Appendix B.
Figure 3: Comparison of the defensive capabilities of diffusion-based and autoregressive LLMs across three jailbreak benchmarks: (a) under the AIM attack (to avoid missing bars due to zero values, all ASR-k scores are uniformly offset by +5%), and (b) under the PAIR attack. Additional experimental results can be found in Figure 8 and Figure 9 of Appendix C.

### v2

@@ 1 Introduction
Diffusion-based language models (dLLMs) (Ye et al., 2025; Nie et al., 2025b; Yang et al., 2025b) have recently emerged as a promising complementary paradigm to traditional autoregressive LLMs (Achiam et al., 2023; Yang et al., 2024a; Wen et al., 2025a; b; 2026; Team et al., 2026). Unlike sequential generation, dLLMs support parallel decoding of masked tokens and leverage bidirectional context modeling, enabling theoretically faster inference and more holistic understanding of input prompts (Yu et al., 2025). These advantages have led to impressive performance and efficiency in tasks such as code generation (Labs et al., 2025; Gong et al., 2025), complex reasoning (Zhu et al., 2025), and text infilling (Li et al., 2025). Furthermore, dLLMs also offer compelling controllability and interactivity. Specifically, users can flexibly insert masked tokens at arbitrary positions in the instruction or generated content, allowing for precise, context-aware editing or rewriting, formatted generation, and structured information extraction as shown in Figure 1.
@@
Although the efficiency and interactivity enabled by parallel decoding and bidirectional context modeling highlight the great potential and promising applications of dLLMs, they may also expose new vulnerabilities in safety. Specifically, dLLMs depart from the left‑to‑right, causal generation paradigm (Achiam et al., 2023; Touvron et al., 2023) and instead leverage bidirectional context modeling. At each denoising step, the dLLMs can “see” the entire token sequence around every [MASK] and fill in tokens that best preserve overall coherence and fluency. While this design yields remarkable flexibility, inserting or regenerating arbitrary spans without retraining, it also creates a blind spot in current safety alignment techniques. By interleaving sets of [MASK] tokens after a vanilla malicious prompt, as shown in Figure 2, a dLLM is coerced into generating harmful instructions purely to maintain contextual consistency.
@@
Moreover, in contrast to autoregressive LLMs, which generate tokens sequentially and can perform on-the-fly rejection of unsafe continuations, dLLMs decode masked tokens in parallel at each step, substantially limiting the model’s ability to conduct dynamic risk assessment or intervene during generation (e.g., rejection sampling for tokens corresponding to harmful content). Consequently, defenses designed for left‑to‑right models break down, opening the door to powerful new jailbreak attacks.
@@
Motivated by these findings, we also take an initial step toward architecture-aware safety alignment for dLLMs. We propose a refusal-aware denoising alignment strategy that trains the model to emit a standardized refusal when confronted with interleaved mask-text jailbreak prompts, rather than completing harmful spans. Please refer to Appendix A for methodology and results.
Our main contributions are summarized as follows:
@@ 2 Related Works
Diffusion Large Language Models. Diffusion Models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have significantly advanced the field of generative modeling, particularly in continuous domains such as images (Rombach et al., 2022; Peebles and Xie, 2023). However, extending these models to discrete data like text introduces distinct challenges due to the inherent discreteness of language. A promising direction in this space is the development of Masked Diffusion Models (MDMs) (Austin et al., 2021; Lou et al., 2023; Shi et al., 2024; Nie et al., 2025a; b; Hoogeboom et al., 2021; Campbell et al., 2022), which generate text by iteratively predicting masked tokens conditioned on their surrounding context. This approach has emerged as a compelling alternative to the traditional autoregressive framework in large language models (LLMs), opening new avenues for text generation. Noteworthy instances of MDMs include LLaDA (Nie et al., 2025b), an 8-billion-parameter model trained from scratch with a bidirectional Transformer architecture, and Dream (Ye et al., 2025), which builds upon pre-trained autoregressive model (ARM) weights. Both models achieve performance comparable to similarly sized ARMs such as LLaMA3 8B (Dubey et al., 2024). The bidirectional nature of these models offers potential advantages over ARMs, including mitigating issues like the reversal curse (Berglund et al., 2023), thus positioning diffusion-based methods as a competitive alternative for next-generation foundation language models.
@@
Jailbreak Attacks and Defenses. Recent studies reveal diverse jailbreak attacks on LLMs by treating them as either computational systems or cooperative agents (Ren et al., 2024; Chen et al., 2024c). Search-based methods like GCG (Zou et al., 2023), AutoDAN (Liu et al., 2023), and PAIR (Chao et al., 2025) use optimization or genetic algorithms to generate adversarial prompts, while side-channel attacks exploit low-resource languages to evade safety checks (Deng et al., 2023). Other techniques target LLMs’ weaknesses in reasoning and symbolic understanding, including scenario nesting (Ding et al., 2023), prompt decomposition (Li et al., 2024), and ASCII obfuscation (Jiang et al., 2024). Additionally, some attacks anthropomorphize LLMs, inducing harmful outputs through persuasion or cognitive overload (Li et al., 2023; Zeng et al., 2024; Xu et al., 2023). To mitigate these threats, defenses fall into four main categories: (1) filter-based detection via perplexity or external classifiers (Jain et al., 2023; Phute et al., 2023; Chen et al., 2024b; ); (2) input modification through permutation or paraphrasing (Robey et al., 2023); (3) prompt-based reminders to reinforce ethical behavior (Xie et al., 2023); and (4) optimization-based approaches such as robust prompt design or safe alignment (Qian et al., 2024; Zhou et al., 2024; Xu et al., 2024; Lu et al., 2025). However, these methods are predominantly developed for autoregressive LLMs. The jailbreak safety of diffusion-based LLMs remains largely unexplored, leaving an important open problem.
@@ 3.1 Preliminary
Let $\mathcal{T}$ be the token vocabulary and $\texttt{[MASK]}\in\mathcal{T}$ the special mask token. Given a prompt $\mathbf{c}=(c_{1},\dots,c_{M})$, the model generates a response $\mathbf{y}=(y_{1},\dots,y_{L})$ through $K$ discrete denoising steps, indexed by $k=K$ down to 0. Let $\mathbf{y}^{(k)}\in\mathcal{T}^{L}$ denote the intermediate state at step $k$, starting from a fully masked sequence:
@@
$P_{\phi}(\mathbf{y}|\mathbf{c},\mathbf{y}^{(k)})=f_{\phi}(\mathbf{c},\mathbf{y}^{(k)};\phi),$ (2)
where $\phi$ represents the model parameters.
@@
Bidirectional Masked Generation. The bidirectional modeling capability and non-autoregressive generation mechanism of dLLMs enable flexible insertion of mask tokens at arbitrary positions in existing text. To accommodate this, unlike standard generation which starts from a fully masked sequence (Eq. 1), the initial state $\mathbf{y}^{(K)}$ is derived by remasking arbitrary spans of an existing sequence for flexible re-generation, resulting in a mix of fixed text tokens (constraints) and mask tokens, where any token $y_{i}\in\mathcal{T}$. The model then performs contextual infilling by iteratively denoising the masked spans using the same transition function $S$ defined in Eq. 4, while keeping the unmasked text tokens fixed. This capability unlocks promising application prospects and enables flexible user interactivity beyond the constraints of traditional autoregressive models, facilitating:
  • Targeted regeneration by masking unsatisfactory spans $\mathbf{y}_{[i:j]}$.
@@ 3.2 DiJA: Diffusion-based LLMs Jailbreak Attack
3.2.1 Problem Formulation
We begin by constructing the corresponding interleaved mask-text jailbreak prompt based on the vanilla harmful prompt (e.g., harmful behaviors from Harmbench (Mazeika et al., 2024)). Let $\mathbf{a}=(a_{1},...,a_{R})$ be a harmful prompt and $\mathbf{m}=(\texttt{[MASK]},...,\texttt{[MASK]})_{Q}$ be $Q$ consecutive masks. An interleaved mask-text jailbreak prompt can be constructed:
@@ 3.2.1 Problem Formulation
where $\oplus$ denotes concatenation, $\otimes$ interleaving, and $\mathbf{w}$ benign separator text. It is worth noting that our constructed prompt does not obscure or remove any of the hazardous content present in the vanilla harmful prompt. This interleaved mask-text prompt construction enables forced generation at specific masked positions, which fundamentally bypasses alignment safeguards in dLLMs. Formally, given an interleaved mask-text prompt $\mathbf{p}_{\text{i}}$, the model’s output distribution factorizes as:
$P_{\phi}(\mathbf{y}|\mathbf{p}_{\text{i}})=\prod_{t\in\mathcal{M}}P_{\phi}(y_{t}|{\mathbf{p}_{\text{i}}\setminus t})\cdot\prod_{t\notin\mathcal{M}}\delta(y_{t}=p_{t}),$ (6)
@@
1:Vanilla harmful prompt $\mathbf{a}=(a_{1},\dots,a_{R})$ $\triangleright$ Source of harmful intent (seed)
@@
3:Examples of interleaved text-mask prompts $\mathcal{E}=\{(\mathbf{a}^{(i)},\mathbf{p}^{(i)}_{\text{i}})\}_{i=1}^{K}$
4:Attacker LLM $\mathcal{L}$; Target Victim dLLM $\mathcal{D}$
@@
6:// Stage 1: Prompt Transformation
@@
8:Provide few-shot examples of interleaved prompts and vanilla harmful prompt $\mathbf{a}$ to $\mathcal{L}$
@@
11:// Stage 2: Masked Decoding (Attack)
@@
14:  Sample token: $y_{t}\sim P_{\phi}(y_{t}\mid\mathbf{p}_{\text{i}}\setminus t)$ $\triangleright$ The decoding of [MASK] is performed in parallel
@@
3.2.2 Method Design
We leverage a language model (e.g., Qwen2.5-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct or GPT-4o) to automatically construct interleaved mask-text jailbreak prompts via in-context learning. The in-context learning template can be found in Appendix C.5. To ensure the generalization and effectiveness of DiJA in jailbreak attacking, we introduce three strategies into the in-context learning process, aiming to enhance the diversity and coherence of the constructed interleaved mask-text jailbreak prompts.
Prompt Diversification. To ensure the diversity of interleaved mask-text jailbreak prompts, it is essential to first guarantee the diversity of the underlying vanilla jailbreak prompts from which they are constructed. We manually curate a small yet diverse set of harmful examples as few-shot demonstrations for in-context learning. These examples span a variety of harmful prompt forms (e.g., step-by-step guides, Q&A, lists, markdowns, dialogues, emails) and harmful content (e.g., malware generation, phishing schemes, hate speech, illegal drug recipes, violence instructions ), ensuring robustness against surface-level prompt variations. We further inject stylistic perturbations (e.g., tone, formality, verbosity) to simulate realistic adversarial scenarios and prevent overfitting.
Masking Pattern Selection. Building on the diversified vanilla prompts, we apply a range of masking strategies to further enhance the diversity of masking patterns. These include: Block-wise masking, which masks entire spans to simulate redacted instructions and elicit long, coherent generations; Fine-grained masking, which selectively hides key tokens (e.g., verbs or entities) while preserving structure; and Progressive masking, which incrementally masks critical information across multi-step instructions to amplify intent. Each strategy balances contextual anchoring with generative freedom, allowing fine-grained control over dLLM behavior and broadening attack coverage. Illustrative examples are provided in Table 15 (Appendix C.5).
Benign Separator Insertion. After ensuring diversity in content and structure of the vanilla prompts as well as in the masking patterns, a crucial step lies in preserving the fluency and coherence of the final interleaved mask-text prompts. This involves carefully aligning the vanilla prompt segments with the masked tokens to maximize the effectiveness of the attack. Thus, we insert short, harmless snippets drawn from a curated phrase pool or generated via controlled prompts. These separators are stylistically consistent (e.g., factual, instructive, narrative), semantically neutral, and capped at ten words. They serve two key purposes: (i) preserving fluency and structural coherence, and (ii) anchoring context to guide dLLMs toward harmful completions. Importantly, the separators are context-sensitive, adapted to the rhetorical style of the original harmful prompt (e.g., procedural, persuasive, or conversational), to ensure seamless integration and stealth. This alignment helps model treat masked spans as natural continuations, improving attack success without sacrificing realism.
@@ 4.1 Experiment Settings
Implementation Details. To evaluate the effectiveness of our proposed automated jailbreak attack pipeline and uncover critical security vulnerabilities in existing diffusion-based LLMs (dLLMs), we conduct experiments on representative dLLMs, including the LLaDA family (Nie et al., 2025b), Dream family (Ye et al., 2025), and MMaDA family (Yang et al., 2025a), across multiple recognized jailbreak benchmarks (Mazeika et al., 2024; Chao et al., 2024a; Souly et al., 2024). We experimented with two LLMs for constructing and refining interleaved mask-text jailbreak prompts in DiJA: Qwen2.5-7B-Instruct (Yang et al., 2024a) (denoted as DiJA) and GPT-4o (Hurst et al., 2024) (denoted as DiJA*), with results reported in Tables 12, and 3, respectively. For more details on the victim models, benchmarks, and baselines, please refer to the Appendix C.
Evaluation Metrics. Building on prior works (Liu et al., 2023; Chao et al., 2023; Ding et al., 2023; Dong et al., 2024; Chen et al., 2024d), we evaluate using GPT-judged Harmful Score (HS) and Attack Success Rate (ASR), including keyword-based ASR (ASR-k) and evaluator-based ASR (ASR-e). GPT-4o rates victim model responses from 1 (refusal or harmless) to 5 (highly harmful or relevant), HS=5 denotes a successful jailbreak. We use the same judging prompt as in previous studies (Qi et al., 2023). For more details on the evaluation metrics, please refer to the Appendix D.
Figure 3: Comparison of the defensive capabilities of diffusion-based and autoregressive LLMs across three jailbreak benchmarks: (a) under the AIM attack (to avoid missing bars due to zero values, all ASR-k scores are uniformly offset by +5%), and (b) under the PAIR attack. Additional experimental results can be found in Figure 8 and Figure 9 of Appendix B.2.